Write a quadratic equation such that the sum of its roots is 6 and the product of its roots is -27
Answer by ankor@dixie-net.com(22740)

Write a quadratic equation such that the sum of its roots is 6 and the product of its roots is -27
Let x = a, and x = b
a + b = 6
b = -a + 6
ab = -27
replace b with (-a+6)
a(-a+6) = -27
-a^2 + 6a + 27 = 0
a^2 - 6a - 27 = 0
(a-9)(a+3) = 0
a = 9, then b = -3
a = -3 then b = 9
:
the quadratic equation: y = x^2 - 6x - 27